Show that, "Three unbalanced phasors of a 3- system can be resolved into three balanced system of phasors."


1
Expert's answer
2021-08-19T08:27:02-0400


Consider an unbalanced phasor has shown in (a). The positive, negative, and zero sequence are shown in (b), (c), and (d) respectively.

Let's define the a-operator

"a=e^{j120\degree}=-\dfrac{1}{2}+j\dfrac{\sqrt{3}}{2}\\
a^2=e^{j240\degree}=-\dfrac{1}{2}-j\dfrac{\sqrt{3}}{2}\\"

from (b)

"V_{b1}=a^2V_{a1}\\
V_{c1}=aV_{a1}\\"

from (c)

"V_{b2}=aV_{a2}\\
V_{c2}=a^2V_{a2}"

from (d)

"V_{a0}=V_{b0}=V_{c0}"



"V_{a}=V_{a0}+V_{a1}+V_{a2}\\
V_{b}=V_{b0}+V_{b1}+V_{b2}\\
V_{b}=V_{a0}+a^2V_{a1}+aV_{a2}\\
V_{c}=V_{c0}+V_{c1}+V_{c2}\\
V_{c}=V_{a0}+aV_{a1}+a^2V_{a2}\\"